# Benchmarking LLMs for Translating Classical Chinese Poetry: Evaluating Adequacy, Fluency, and Elegance

### v1

@@ 4.2 The Second Workflow
3.Extractor. The goal of the Extractor is to extract the final translation results generated by the Voter. As an agent of the LLM, the Extractor filters out noise from the content generated by the Voter based on the source input, and outputs the final translation results. Specific prompts are displayed in Appendix B.3.
5 Evaluation Methods
5.1 Evaluation Criteria
The translation of classical poetry requires not only artistic expression but also an understanding of the cultural background, making it nearly impossible to produce a single, universally correct translation. Therefore, multiple translations of the same classical poem can coexist, each offering a unique perspective and interpretation. Following this line of thought, we evaluate the quality of classical poetry translations through three aspects: “Adequate," “Fluent," and “Elegant." The specific evaluation methods are as follows:
Figure 4: Examples of different evaluation metrics.
  1. 1.
    Adequate
    1. (a)
  2. 2.
    Fluent
    1. (a)
      Beauty of Sound (BS)↑: The beauty of sound in Chinese classical poetry is primarily reflected in its rhyme and meter. This standard examines whether the translation achieves harmonious sound, precise wording, adherence to strict metrical rules, and a rhythm that is both smooth and dynamic. As shown in Figure 4(b).
    2. (b)
      Beauty of Form (BF)↑: Chinese classical poetry emphasizes balanced structure, with common forms such as "four-character quatrains," "seven-character quatrains," and "seven-character regulated verse," each exhibiting distinct aesthetic characteristics of Chinese poetry. This standard evaluates whether the translation maintains consistency with the source poem’s structure, including the alignment of line numbers and balanced phrasing. As shown in Figure 4(c).
  3. 3.
    Elegant
    1. (a)
      Beauty of Meaning (BM)↑: Chinese classical poetry uses concise and precise language to create vivid imagery and a rich atmosphere for readers. This standard assesses the depth and richness of the translation, focusing on the effectiveness of conveying themes, emotions, and messages. As shown in Figure 4(a).
5.2 LLM-based Classical Poetry Metric
In this section, we aim to establish a method for evaluating the translation of classical Chinese poetry based on LLMs assessments. Drawing from current QE (Quality Estimation) research, we have developed an approach to utilize LLM in determining the quality of classical Chinese poetry translations. As shown in Figure 5, we designed a 1-5 scoring prompt to help LLM focus on the translation quality in terms of Beauty of Sound (LLM-BS). In the evaluation criteria, a score of 1 reflects a poor translation of classical Chinese poetry, a score of 3 indicates a basic but imperfect translation result and a score of 5 represents an excellent and accurate translation in this aspect. Subsequently, test data is provided to the LLM, and the LLM is asked to generate an assessment with scores only. Additionally, we calculate the average scores of BF, BF, and BM (LLM-Avg) to evaluate the overall translation performance. The prompt details for Beauty of Form (LLM-BF) and Beauty of Meaning (LLM-BM) are provided in Appendix F and G.
@@ 5.2 LLM-based Classical Poetry Metric
Evaluate the beauty of sound in the given Chinese translation of classical poetry. Focus on whether the translation achieves harmonious sound, precise wording, adherence to strict metrical rules, and a rhythm
@@
5.3 Traditional Automatic Evaluation
In evaluating our translation methodology, we initially employ COMET [^1] https://huggingface.co/Unbabel/wmt22-comet-da Rei et al. (2022) and BLEURT [^2] https://github.com/lucadiliello/bleurt-pytorch Sellam et al. (2020b) as automatic metrics, aligning with the established standards in LLM-based translation literature He et al. (2023); Huang et al. (2024). For traditional translation evaluation, we use BLEU  [^3] https://github.com/mjpost/sacrebleu Papineni et al. (2002).
6 Experiment Setup
6.1 Comparing Systems
In our evaluation, RAT is compared with a range of translation methods, including Zero-shot Wei et al. (2022), 5-shot Brown et al. (2020); Hendy et al. (2023), Rerank Moslem et al. (2023a), Refine Chen et al. (2023), MAD Liang et al. (2023), and MAPS He et al. (2023). To validate its generalizability, we utilize three LLMs, which include closed-source models such as ChatGPTOuyang et al. (2022) and GPT-4Achiam et al. (2023)  [^4] The ChatGPT and GPT-4 models used in this work are accessed through the gpt-3.5-turbo and gpt-4 APIs, respectively., as well as open-source models like Alpaca-7B Taori et al. (2023)  [^5] https://huggingface.co/tatsu-lab/alpaca-7b-wdiff/tree/main and Vicuna-7B Chiang et al. (2023)  [^6] https://huggingface.co/lmsys/vicuna-7b-v1.5. The following content will provide detailed descriptions of these comparisons:
  • Rerank Moslem et al. (2023a) was conducted with the identical prompt as the baseline, employing a temperature of 0.3 (Moslem et al., 2023b). Three random samples were generated and combined with the baseline to yield four candidates. The optimal candidate was chosen through GPT-4.
  • Renfie Chen et al. (2023) first requests a translation from ChatGPT, then provides the source text and translation results, and obtains a refined translation through multiple rounds of modifications by mimicking the human correction process.
  • MAD Liang et al. (2023) enhance the capabilities of large language models (LLMs) by encouraging divergent thinking. In this method, multiple agents engage in a debate, while an agent oversees the process to derive a final solution.
  • Dual-ReflectChen et al. (2024b) provide supervisory signals for large models to reflect on translation results through dual learning, thereby iteratively improving translation performance (the maximum number of iterations is set to 5).
  • RAT,is proposed method in this work.
7 Experimental Results
7.1 Can LLM evaluate Chinese classical Poetry ?
We first randomly select N chapter-level Chinese classical poetry pieces from each source language dataset and translate them using the RAT method. We then compare the evaluation results of the large language model with human-annotated results to calculate the Pearson correlation coefficient Pearson (1920), Spearman correlation coefficient Spearman (1961), and Kendall correlation coefficient Kendall (1948) to determine the level of consistency.
Metric Pearson’s $r$ Spearman’s $\rho$ Kendall’s $\tau$
BLEU -0.23 -0.18 -0.12
BLEU-1 0.05 0.08 0.05
BLEURT 0.14 0.16 0.11
COMET 0.13 0.18 0.11
LLM-BM 0.85 0.81 0.85
LLM-BF 0.71 0.75 0.70
LLM-BS 0.73 0.73 0.76
LLM-AVG 0.77 0.73 0.75
The results in Table N indicate that large language models can serve as effective tools for evaluating the translation quality of idioms across different language pairs. In contrast, BLEU, COMET, and BLEURT were found to be less consistent with human evaluations, highlighting the advantages of our proposed evaluation method in the translation of classical poetry.
7.2 Main Results
Methods Discourse-Level Poetry Transaltion
COMET BLEURT LLM-BM LLM-BS LLM-BF LLM-Avg BLEU BLEU-1 BLEU-2 BLEU-3 BLEU-4
GPT-4 60.3 43.0 4.0 3.7 3.6 3.8 4.0 22.1 7.8 3.3 1.7
ChatGPT 61.1 42.4 3.3 3.2 2.9 3.1 4.3 23.4 8.7 3.1 1.8
   +5shot 61.0 42.5 3.5 3.3 3.3 3.4 3.7 22.0 7.7 3.2 1.6
   +rerank 61.0 42.5 3.7 3.7 3.9 3.8 3.7 22.5 8.0 3.4 1.7
   +MAD 59.9 42.3 3.7 3.6 3.8 3.8 4.3 23.2 8.8 3.7 1.8
   +Dual-Reflect 58.2 40.9 3.8 3.8 3.9 3.9 3.8 20.5 7.5 3.2 1.6
   +EAPMT 61.1 42.9 3.8 3.7 3.8 3.7 3.7 21.6 7.5 3.1 1.5
   +RAT 62.7 43.9 4.1 3.9 3.9 4.0 4.6 23.9 9.8 3.9 2.2
Llama3-8B 54.3 37.4 2.7 2.6 2.4 2.5 1.7 17.4 6.1 3.5 1.3
   +5shot 54.5 37.6 2.9 2.8 2.6 2.7 1.8 17.4 6.2 3.4 1.3
   +Rerank 54.8 38.1 3.0 3.3 3.5 3.2 2.0 17.9 6.6 3.6 1.5
   +RAT 55.6 38.4 3.4 3.3 3.6 3.4 2.2 18.2 7.0 3.9 1.8
Vicuna-7B 52.2 26.4 2.4 2.4 1.8 2.2 0.3 16.5 4.7 3.4 1.0
   +5shot 52.4 26.1 2.5 2.6 2.3 3.6 1.0 17.1 4.3 3.6 1.3
   +Rerank 52.8 26.3 3.0 2.6 3.3 2.8 1.2 17.5 5.0 3.7 1.6
   +RAT 60.1 26.9 3.0 2.5 3.3 2.9 1.8 17.6 5.3 3.9 1.9
We compare various different LLM-based methods on the Classical Chinese Poetry Benchmark with RAT. The analysis of the experimental results is as follows:
  • The effectiveness of different LLMs. Among all comparative methods, closed-source models perform better on this task than open-source models, possibly implying that closed-source models benefit from richer pre-training data, thus enabling higher-quality translations. This also suggests that the Chinese Classical Poetry translation task is more challenging.
  • The effectiveness of retrieved knowledge. For EAPMT, which relies on LLMs’ self-generated knowledge, the RAT method, which uses retrieval-based knowledge, provides more accurate information and further improves translation quality. This suggests that translating Chinese Classical Poetry demands stricter knowledge. However, the results from RAT still do not reach the level of human translation, indicating significant room for improvement in this task.
7.3 Evaluation of Adequacy
To evaluate the translation performance of LLMs in terms of Adequacy, we employed a constructed dataset of 500 Classical Chinese Sentence-Level Translations to assess various translation methods. This experiment follows the method of Liang et al., 2023, assessing translation results from two main dimensions: manual evaluation of ambiguous translation accuracy (see Appendix A for details) and direct evaluation of translation quality using the LLM-BM.
The experimental results are shown in Table 2. In terms of ambiguity resolution accuracy, RAT performed the best, indicating that retrieving accurate information helps to better eliminate ambiguities in translation tasks. In the human evaluation, RAT received the highest score, further demonstrating that this method achieves excellent translation quality.
Methods LLM-BM ACC
GPT-4 3.5 57.4
ChatGPT
   +Zero-Shot 2.9 51.8
   +Rerank 3.0 56.4
   +Dual-Reflect 3.3 54.3
   +MAD 3.5 56.1
   +RAT 3.8 57.8
Table 2: The human-annotated results of the Adequacy in Sentence-Level Commonsense MT benchmark Translation.
8 Analysis
8.1 LLM-based Metric Consistency
In this experiment, we evaluated whether the proposed LLM-based metrics, including LLM-BS, LLM-B, LLM-BM, and LLM-AVG, could reflect the evaluation characteristics of Beauty of Sound, Beauty of Form, Beauty of Meaning, and overall translation quality, while also demonstrating variability. We conducted pairwise correlation tests between human evaluation results and LLM-based results, employing the Pearson correlation coefficient, Spearman correlation coefficient, and Kendall correlation coefficient. The experimental results are illustrated in Figure N.
8.2 Impact of Different Knowledge on Translation Performance
The RAT method utilized knowledge from the Classical Chinese Poetry Knowledge Base for translation. To investigate which specific knowledge is more effective in translating Classical Chinese Poems within knowledge-based translation methods, this experiment employed a modified RAT method that used only one type of retrieved knowledge at a time and removed the Voter module (Figure 4). The experimental results are presented in Figure 7.
Figure 7: Experiment on the Impact of Different Knowledge of Classical Chinese Poetry on Translation
On the one hand, retrieval-based translation methods showed overall performance improvement, indicating that knowledge plays a significant role in the task of translating Classical Chinese Poetry. On the other hand, the results revealed that knowledge of modern Chinese translation contributed the most to performance enhancement. This suggests that there are differences in form, semantics, and rhythm between Classical Chinese Poems and modern Chinese, further demonstrating the considerable challenges of directly translating Classical Chinese Poetry.
To explore the translation difficulty of different types of poems (Tang, Song, and Yuan) in the collected 608 Chinese Classical Poems, we employed the RAT method to generate translation results for each poetry type. These results were then evaluated using LLM-BF, LLM-BM, LLM-BS, and LLM-AVG. The experimental results are shown in Figure 8.
From the experimental results, we can observe that the translation trends are consistent across different types. Firstly, the translation of Tang poetry is relatively easier compared to other types, likely due to its stricter format and fewer words. Secondly, the lower scores of LLM-BF and LLM-BS indicate that the translation of poetic structure and rhythm remains a significant challenge for LLMs. Lastly, the generally higher results of LLM-BM suggest that retrieval-based methods can produce more adequate translations.
9 Conclusion
Our research reveals the challenges LLMs face in translating classical Chinese poetry, especially in cultural knowledge, fluency, and elegance. We introduced the Retrieval-Augmented Translation (RAT) method and a GPT-4-based evaluation metric to improve translation quality. This first comprehensive evaluation exposes LLMs’ limitations and suggests areas for improvement, aiming to inspire better culturally informed translations.
10 Acknowledgements
Appendix A Human Evaluations
In this section, we conduct a human evaluation to measure translation quality. We assess ambiguity resolution. Four native English speakers were invited to participate. In the Adequacy sentence-level task, the four experts scored each sentence for ambiguity resolution against the reference, awarding 1 point for resolved and 0 points for unresolved.
A.1 Detailed prompt for Selector

### v4

@@ 5.2 The Second Workflow
Extractor. The goal of the Extractor is to extract the final translation results generated by the Voter. As an agent of the LLM, the Extractor filters out noise from the content generated by the Voter based on the source input and outputs the final translation results. Specific prompts are displayed in Appendix B.4.
6 Experiment Setup
6.1 Comparing Systems
In our evaluation, RAT is compared with a range of translation methods, including Zero-shot Wei et al. (2022), 5-shot Hendy et al. (2023), Rerank Moslem et al. (2023a), Refine Chen et al. (2023), MAD Liang et al. (2023), EAPMT Wang et al. (2024), and Dual-Reflect Chen et al. (2024a). To validate its generalizability, we utilize three LLMs, which include closed-source models such as ChatGPTOuyang et al. (2022) and GPT-4Achiam et al. (2023)  [^4] The ChatGPT and GPT-4 models used in this work are accessed through the gpt-3.5-turbo and gpt-4 APIs, respectively., as well as open-source models like Llama3-8B Dubey et al. (2024)  [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B and Vicuna-7B Chiang et al. (2023)  [^6] https://huggingface.co/lmsys/vicuna-7b-v1.5. Details on comparative methods are in Appendix B.5.
6.2 Evaluation Metrics
LLM-based Automatic Evaluation. We propose an automatic evaluation method for translation based on LLMs in the section 4.
Traditional Automatic Evaluation. We initially employ COMET [^7] https://huggingface.co/Unbabel/wmt22-comet-da Rei et al. (2022) and BLEURT [^8] https://github.com/lucadiliello/bleurt-pytorch Sellam et al. (2020) as automatic metrics, aligning with the established standards in LLM-based translation literature He et al. (2023); Huang et al. (2024). For traditional translation evaluation, we use BLEU  [^9] https://github.com/mjpost/sacrebleu Papineni et al. (2002).
7 Experimental Results
Methods Discourse-Level Poetry Translation
COMET $\uparrow$ BLEURT $\uparrow$ LLM-BM $\uparrow$ LLM-BS $\uparrow$ LLM-BF $\uparrow$ LLM-Avg $\uparrow$ BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ BLEU-3 $\uparrow$ BLEU-4 $\uparrow$
GPT-4 60.3 43.0 4.0 3.7 3.6 3.8 22.1 7.8 3.3 1.7
ChatGPT 61.1 42.4 3.3 3.2 2.9 3.1 23.4 8.7 3.1 1.8
   +5shot 61.0 42.5 3.5 3.3 3.3 3.4 22.0 7.7 3.2 1.6
   +Rerank 61.0 42.5 3.7 3.7 3.9 3.8 22.5 8.0 3.4 1.7
   +MAD 59.9 42.3 3.7 3.6 3.8 3.7 23.2 8.8 3.7 1.8
   +Dual-Reflect 58.2 40.9 3.8 3.8 3.9 3.8 20.5 7.5 3.2 1.6
   +EAPMT 61.1 42.9 3.8 3.7 3.8 3.7 21.6 7.5 3.1 1.5
   +RAT 62.7 43.9 4.1 3.9 3.9 4.0 23.9 9.8 3.9 2.2
Vicuna-7B 52.2 26.4 2.4 2.4 1.8 2.2 16.5 4.7 3.4 1.0
   +5shot 52.4 26.1 2.5 2.6 2.3 2.4 17.1 4.3 3.6 1.3
   +Rerank 52.8 26.3 3.0 2.6 3.3 2.8 17.5 5.0 3.7 1.6
   +RAT 60.1 26.9 3.0 2.5 3.3 2.9 17.6 5.3 3.9 1.9
Llama3-8B 54.3 37.4 2.7 2.6 2.4 2.5 17.4 6.1 3.5 1.3
   +5shot 54.5 37.6 2.9 2.8 2.6 2.7 17.4 6.2 3.4 1.3
   +Rerank 54.8 38.1 3.0 3.3 3.5 3.2 17.9 6.6 3.6 1.5
   +RAT 55.6 38.4 3.4 3.3 3.6 3.4 18.2 7.0 3.9 1.8
Qwen-72B 60.9 43.5 3.4 3.4 3.3 3.3 22.1 7.1 3.0 2.0
   +5shot 60.4 43.8 3.6 3.7 3.4 3.5 21.5 7.2 2.9 1.5
   +Rerank 59.8 43.2 3.0 3.3 3.5 3.2 20.6 6.7 2.7 1.3
   +RAT 61.7 43.5 3.7 3.6 3.6 3.6 22.9 8.0 2.9 2.0
Table 2: The main results from the PoetMT benchmark are presented. The bold indicates the highest scores. The bolded results indicate the highest statistically significant scores (p-value $<0.05$ in the paired t-test against all compared methods).
7.1 Can LLM evaluate Classical Poetry ?
We first translate randomly selected 100 discourse-level data from the PoetMT benchmark by the RAT method. Inspired by Li et al., 2023, we then scored the translations according to the citeria outlined in Figures 7, 8, and 9. We compare the different evaluation results of the different automatic methods with human-annotated results to calculate the Pearson correlation coefficient Pearson (1920), Spearman correlation coefficient Spearman (1961), and Kendall correlation coefficient Kendall (1948) to determine the level of consistency.
Metric Pearson’s $r$uparrow Spearman’s $\rho$uparrow Kendall’s $\tau$uparrow
Traditional Automatic Evaluation
BLEU -0.23 -0.18 -0.12
BLEU-1 0.05 0.08 0.05
BLEURT 0.14 0.16 0.11
COMET 0.13 0.18 0.11
Qwen-72B-based Automatic Evaluation
LLM-BM 0.63 0.59 0.61
LLM-BF 0.53 0.55 0.50
LLM-BS 0.54 0.53 0.55
LLM-AVG 0.57 0.53 0.54
GPT-4-based Automatic Evaluation
LLM-BM 0.85 0.81 0.85
LLM-BF 0.71 0.75 0.70
LLM-BS 0.73 0.73 0.76
LLM-AVG 0.77 0.73 0.75
Table 3: correlation metrics between human and BLEU, BLEU-1, COMET, BLEURT, LLM-BM, LLM-BF, LLM-BS or LLM-AVG evaluation on our PoetMT.
The results in Table 3 indicate that large language models can serve as effective tools for evaluating the translation quality of classical Chinese poetry. In contrast, BLEU, COMET, and BLEURT were found to be less consistent with human evaluations, highlighting the advantages of our proposed evaluation method in the translation of classical poetry.
7.2 Main Results
We compare various different LLM-based methods on the PoetMT benchmark with RAT. The results are shown in Table 2. The analysis of the experimental results is as follows:
Performance Variations Among Different Types of LLMs. Among all comparative methods, closed-source models perform better on this task than open-source models, possibly implying that closed-source models benefit from richer pre-training data, thus enabling higher-quality translations. This also suggests that the PoetMT task is more challenging.
The effectiveness of retrieved knowledge. For methods that rely on LLMs’ self-generated knowledge (such as the EAPMT method), the RAT approach, which utilizes retrieval-based knowledge, provides more accurate information and thus enhances translation quality. This suggests that more accurate relevant knowledge can better assist in the PoetMT task.
7.3 Evaluation of Adequacy
To evaluate the translation adequacy of LLMs, we used a dataset of 758 Classical Chinese sentence-level translations. Following Liang et al., 2023 and Chen et al., 2024b, we evaluated translations through manual adequacy assessment (Appendix A.1), the LLM-BM score (GPT-4), and the human BM score (Appendix A.2).
Results (Table 10) show that RAT achieved the best adequacy scores. This suggests that retrieving accurate information improves adequacy. RAT also received the highest LLM-BM score, confirming that it better captures the themes, emotions, and messages of the original poems.
Methods LLM-BM $\uparrow$ Human-BM $\uparrow$ ACC $\uparrow$
GPT-4 3.9 3.6 69.1
ChatGPT
   +Zero-Shot 3.2 3.2 60.5
   +Rerank 3.2 3.3 64.4
   +Dual-Reflect 3.7 3.6 66.4
   +MAD 3.7 3.8 67.3
   +RAT 3.9 3.9 69.9
Vicuna-7B
   +Zero-Shot 2.1 0.8 26.9
   +Rerank 2.3 1.2 31.7
   +Dual-Reflect 2.0 1.0 33.0
   +MAD 2.2 1.3 67.3
   +RAT 2.5 2.1 43.4
Table 4: The LLM-BM and human-annotated results of the Adequacy in Sentence-Level PoetMT benchmark Translation. The results of Llama3-8B and Qwen-72B are in Appendix C.2.
7.4 Data Validation Experiments
To explore whether PoetMT poems were included in the training data of closed-source LLMs like GPT-4 and ChatGPT (§7.2), we conducted an experiment using 150 poems (50 each from Tang poetry, Song lyrics, and Yuan opera). Following concerns raised by Shi et al. (2024), we prompted GPT-4/ChatGPT with the title and author to generate poems, then evaluated the similarity to human reference using SacreBLEU. As shown in Table 5, the results indicate higher BLEU scores for Chinese and lower scores for English, suggesting fewer English translations in LLM training data.
Type of Poetry Tang Song Yuan
Language Chinese English Chinese English Chinese English
ChatGPT 6.6 0.4 4.4 0.6 1.7 0.4
GPT4 8.1 0.8 7.3 0.9 4.2 0.6
Table 5: BLEU Scores from data validation experiments
7.5 Impact of Different Knowledge on Translation Performance
The RAT method uses the Classical Chinese Poetry Knowledge Base for translation. To find out which type of knowledge is most helpful, this experiment adjusted the RAT method by using only one type of knowledge at a time and removing the Voter module (Figure 4). The results (Figure 5) showed that retrieval-based methods improved overall performance, proving that knowledge is important in translating Classical Chinese Poetry. Among the knowledge types, modern Chinese translation knowledge helped the most, suggesting that differences in style, meaning, and rhythm between Classical and modern Chinese make direct translation challenging.
Figure 5: Experiment on the Impact of Different Knowledge of Classical Chinese Poetry on Translation. The dashed line indicates not using knowledge, but directly translating the result through ChatGPT.
7.6 Ablation Study on Modern Chinese Translations in RAT Framework
In Section 7.5, the Modern Chinese translations in the Knowledge Base within the RAT framework show the greatest impact on output quality. To determine whether the improved results are solely due to these Modern Chinese translations, we conducted an ablation experiment and a case study to analyze this question in greater depth.
COMET $\uparrow$ BLEURT $\uparrow$ LLM-BM $\uparrow$ LLM-BS $\uparrow$ LLM-BF $\uparrow$ LLM-Avg $\uparrow$ BLEU-1 $\uparrow$
ChatGPT-RAT 61.1 42.4 3.3 3.2 2.9 3.1 9.8
ChatGPT-RAT
(only MC)
57.2 38.1 3.1 2.6 2.7 2.8 7.5
Vicuna-RAT 60.1 26.9 3.0 2.5 3.3 2.9 17.6
Vicuna-RAT
(only MC)
53.1 26.9 2.7 2.4 2.5 2.5 19.0
Table 6: Ablation study results comparing the RAT framework with and without Modern Chinese (MC) translations in the Knowledge Base.
In Table 6, the experimental results demonstrate that while Modern Chinese positively influences translation, the RAT method based on multi-knowledge still outperforms it. To further illustrate the limitations of Modern Chinese-based translation, we present three examples in Table 7.6 (additional two examples in the Appendix C.3).
Source: {CJK*}UTF8gbsn红豆生南国,春来发几枝?愿君多采撷,此物最相思
RAT: Red beans grow in the south, sprouting many branches in spring. Pick them often, as they hold deep feelings of longing.
RAT-only Modern Chinese: Red beans grow in the sunny south, sprouting countless new branches every spring. I hope those who are missed will pick more of them, as they best express longing and love.
Reference: Red beans grow in the southern land, In spring, how many branches sprout? I wish you would gather them often, For they most evoke longing thoughts.
Table 7: Comparison of RAT, RAT-only Modern Chinese, and Reference Translations.
The results of the case further support our viewpoint. Translations based on Modern Chinese tend to align more closely with general-domain sentence or passage-level translations and lack the adequacy, fluency, and elegance characteristics.
To analyze the translation difficulty of different types of Classical Chinese poems (Tang, Song, Yuan) from a set of 608 poems, we used the RAT method and evaluated the results with LLM-BF, LLM-BM, LLM-BS, and LLM-AVG (Figure 6).
The results show consistent trends across types. Tang poetry is relatively easier to translate due to its stricter structure and brevity. Lower LLM-BF and LLM-BS scores highlight the challenge of translating poetic structure and rhythm. Higher LLM-BM scores indicate that retrieval-based methods produce more elegant translations.
7.8 Human-centered Error Analysis
To evaluate the RAT method’s effectiveness and limitations, we manually assessed 50 randomly selected poems from the 608 test samples. Using a ChatGPT-based RAT method, translations were scored on a 1-5 scale for semantic adequacy, fluency, and elegance (Figures 7, 8, 9). As shown in Table 8, the results indicate a low proportion of Excellent (5-4) translations and a high proportion of Failed (2-1) ones. This highlights the challenges in PoetMT and the need to further refine the RAT method.
Categories Number of Sentences Rate
Excellent 5 10%
Decent 23 46%
Failed 22 44%
Table 8: Manual evaluation results of 50 RAT translations, categorized by performance.
Based on the results in Table 8, we manually categorized the failed outcomes and provided case examples for clearer illustration in Table 7.8.
Categories Rate Examples: Source/Error Result/Reference
Errors in handling polysemous words 2/22 Source: {CJK*}UTF8gbsn万壑树参天
Error: The trees in your valley scrape the sky
Right: In myriad gorges, trees touch the sky
Lack of cultural context 7/22 Source: {CJK*}UTF8gbsn秦时明月汉时关
Error: The moon still shines on mountain passes as of yore
Right: Under the Qin moon, by the Han frontier
Confusion in long sentence structures 6/22 Source: {CJK*}UTF8gbsn子弟每是个茅草岗沙土窝初生的兔羔儿乍向围场上走
Error: The young gallants are new-born bucks in chase of bunny
Right: Young ones are like rabbits, new to the hunt, Born in a thatch of grass, on sandy ground
Incorrect translation of low-frequency vocabulary 7/22 Source: {CJK*}UTF8gbsn缚虎手
Error: Binding a tiger with bare hands
Right: Barehanded tiger fighting
Table 9: Translation Error Types with Examples.
8 Conclusion
Our research reveals the challenges LLMs face in translating classical Chinese poetry, especially in cultural knowledge, fluency, and elegance. We introduced a GPT-4-based evaluation metric, demonstrating that existing large language models fall short in this task. Additionally, we introduced the RAT method to improve translation quality. This study is the first to evaluate the limitations of large language models in the translation of classical poetry It aims to inspire discussion within the translation community about the potential and future development of LLMs.
9 Limitations
The inherent challenges of translating classical poetry, such as the preservation of rhyme, tone, and aesthetic qualities, remain complex and subjective. Although the proposed GPT-4-based automatic evaluation metric has demonstrated consistency with human evaluation, these subjective dimensions still pose a significant challenge.
Appendix A Human Evaluations
A.1 Human Evaluation for ACC
In this section, we conduct a human evaluation to measure translation quality. We evaluated the adequacy of the translation. Four native English speakers were invited to participate. In the sentence-level adequacy task, the four experts scored each sentence for adequacy against the reference, awarding 1 point for fully adequate and 0 points for inadequate.
A.2 Human Evaluation for BM score
We conducted a human evaluation to measure the BM score, inviting four native English speakers to participate. A sample of 50 sentences was selected from the 758 data points, and the evaluators scored the translations based on the criteria in Figure 9.
Appendix B Detail Prompt
B.1 Detailed prompt for Selector
{mdframed}
[backgroundcolor=purple!10, linecolor=white, linewidth=2pt, roundcorner=10pt]Part-1: Selector: Please identify the knowledge related to the content to translating this classical Chinese poem {text} from the {rag context} knowledge base.
B.2 Detailed prompt for Translator
{mdframed}
B.3 Detailed prompt for Voter
{mdframed}
B.4 Detailed prompt for Extractor
{mdframed}
B.5 Comparative Methods
The following content will provide detailed descriptions of these comparative methods:
  • Rerank Moslem et al. (2023a) was conducted with the identical prompt as the baseline, employing a temperature of 0.3 (Moslem et al., 2023b). Three random samples were generated and combined with the baseline to yield four candidates. The optimal candidate was chosen through GPT4.
  • Refine Chen et al. (2023) first requests a translation from ChatGPT, then provides the source text and translation results, and obtains a refined translation through multiple rounds of modifications by mimicking the human correction process.
  • MAD Liang et al. (2023) enhance the capabilities of LLMs by encouraging divergent thinking. In this method, multiple agents engage in a debate, while an agent oversees the process to derive a final solution.
  • EAPMT Wang et al. (2024) leverages the explanation of monolingual poetry as guidance information to achieve high-quality translations from Chinese poetry to English poetry.
  • Dual-ReflectChen et al. (2024a) provide supervisory signals for large models to reflect on translation results through dual learning, thereby iteratively improving translation performance (the maximum number of iterations is set to 5).
  • RAT is the proposed method in this work.
B.6 Detailed prompt for Beauty of Sound
For evaluation of the beauty of form, the detailed prompt is displayed in Figure 8
@@ B.6 Detailed prompt for Beauty of Sound
Evaluate the beauty of sound in the given Chinese translation of classical poetry. Focus on whether the translation achieves harmonious sound, adherence to strict metrical rules, and a rhythm
@@
B.7 Detailed prompt for Beauty of Form
For evaluation of the beauty of form, the detailed prompt is displayed in Figure 8